# Visual Fixation-Based Retinal Prosthetic Simulation

### v1

@@ 1 Introduction
Fig. 1: (a) Original image. (b) Trace of saccades of the human eye viewing a still image ([14] under CC BY-SA 2.0). (c) Self-attention from a Vision Transformer in DINOv2 [15].
@@ 2.1 Dataset
Fig. 4: A virtual retinal implant with 196 electrodes (14$\times$14) on the retinal axon map [2, 20].
@@ 2.2 Fixation Predictor
We leverage the self-attention [21] of the [CLS] token on the heads of the last layer from an ImageNet pre-trained Vision Transformer (ViT [16]) and regard the attention score as the saliency of the fixation in the saccade mechanism of the human eye [11]. As illustrated in Fig. 1, the self-attention mask highlights regions similar to those traced by saccades. Out of the 256 patches (16$\times$16) from each input image (resized to 224$\times$224 pixels), the top 10% (i.e., int(25.6)=25) most salient fixation patches are preserved based on the sum within each patch of the multi-head attention [21], as shown in Fig. 2.

### v2

@@ 1 Introduction
Fig. 1: (a) Original image. (b) Trace of saccades of the human eye viewing a still image ([14] under $C$ BY-SA 2.0). (c) Self-attention from a Vision Transformer in DINOv2 [15].
@@ 2.1 Dataset
Fig. 4: A virtual retinal implant with 196 electrodes ($\times14$) on the retinal axon map [2, 20].
@@ 2.2 Fixation Predictor
We leverage the self-attention [21] of the [CLS] token on the heads of the last layer from an ImageNet pre-trained Vision Transformer (ViT [16]) and regard the attention score as the saliency of the fixation in the saccade mechanism of the human eye [11]. As illustrated in Fig. 1, the self-attention mask highlights regions similar to those traced by saccades. Out of the 256 patches ($\times16$) from each input image (resized to $\times224$ pixels), the top 10% (i.e., int(25.6)=25) most salient fixation patches are preserved based on the sum within each patch of the multi-head attention [21], as shown in Fig. 2.